you must prepare a 100mL solution of hydrochloric acid with a concentration of 0.500mol/L. you are given the appropriate glassware and a bottle of 2.50mol/L stock hydrochloric acid. What volume of stock solution must you use to prepare the solution?
1
Expert's answer
2020-06-10T07:40:51-0400

Solution.

"C1 \times V1 = C2 \times V2"

"V1 = \frac{C2 \times V2}{C1}"

V1 = 20 ml

Answer:

V1 = 20 ml